OpenStudy (anonymous):

I'LL GIVE MEDAL!!!! The midpoint of UV is (5 -11). The coordinates of one are U(3,5). Find the coordinates of endpoint V. Show your work.

OpenStudy (johnweldon1993):

Well a midpoint is halfway from U to V so \[\large \text{x-coordinate} = \frac{x_1 + x_2}{2}\] \[\large \text{y-coordinate} = \frac{y_1 + y_2}{2}\] Since we know the midpoint is (5,-11) and U = (3,5) we can rewrite these equations as \[\large 5 = \frac{3 + x_2}{2}\] \[\large -11 = \frac{5 + y_2}{2}\] And solve for 'x' and 'y'